# How hydrodynamic interactions alter polymer stretching in turbulence

### v1

@@ V Random gradients emulate turbulence
Before concluding, we check whether the effects of HI on turbulent polymer stretching can be predicted by using random Gaussian velocity gradients, i.e., by replacing the turbulent flow with a synthetic random flow. As discussed in Sec. II.3, the time series of random gradients is constructed so as to have the same Lyapunov exponent and similar temporal correlation properties as the Lagrangian trajectories from the DNS (see Appendix A for details). If the Gaussian gradients work well, then one can study the extensional dynamics of polymer chains without having to perform a DNS. This would facilitate the development of coarse-grained models, like dumbbells with a conformation-dependent drag [14, 16, 11], by allowing for the comparison of polymer models in a fluctuating flow without having to simulate the Navier-Stokes equations. Data-driven modelling approaches [74, 75, 76, 77] which require large amounts of data would particularly benefit from using the random flow model, since one can then perform very long simulations without the need for long Lagrangian trajectories from a DNS.
@@
The PDF of extension predicted using the Gaussian gradients (GG) are compared to the results from the DNS, for twenty-bead chains without HI in Fig. 8(a). The comparison with HI is presented in Fig. 8(b) for $h^{\star}=0.2$ and in Fig. 8(c) for for $h^{\star}=0.49$. The DNS and GG results agree reasonably well even in the presence of HI. Importantly, the differences between the PDFs from the DNS and the GG model are much smaller than the modifications produced by HI [compare the differences in Fig. 8(b) to those in Fig. 4(b)]. This means that we can replace the DNS by the GG random model without compromising our ability to study and understand the effects of HI on polymer stretching in fluctuating flows.
After averaging over the PDFs, the results for the mean extension $\langle R\rangle/R_{\mathrm{eq}}$ from the DNS and the GG random model match well, as shown in Fig. S4 of the supplemental material [69]. We have also found that the effects of HI on the persistence time in stretched and coiled states (discussed in Sec. IV.3) are reproduced by the GG random model.
@@ VI Concluding remarks
Simulations of turbulent polymer solutions are typically based on the elastic dumbbell model, either indirectly via continuum constitutive models like Oldroyd-B or FENE-P [78], or directly via two-way coupled Eulerian-Lagrangian (EL) simulations in which explicit dumbbells are evolved in tandem with the Navier-Stokes equations [79, 80, 4, 47, 81]. While the EL approach is computationally expensive and cannot offer analytical insight, it can be extended naturally to more realistic models for polymers [see, e.g, the recent work by Serafini et al. [82], wherein EL simulations of turbulent pipe flow are performed with bead-spring and bead-rod chains].
@@
Replacing dumbbells with chains, however, greatly increases the computational expense of EL simulations. There must therefore be a good reason to use high-order polymer models. The most obvious characteristic of multibead chains, which is absent in dumbbells, is that elastic relaxation occurs on multiple time-scales. However, this feature alone does not lead to differences in the dynamics of the end-to-end extension. Indeed, the transient and stationary statistics of a Rouse chain in a homogeneous isotropic turbulent flow can be mapped to that of a dumbbell using the Jin-Collins mapping [39]. Our current work has shown that this simple equivalence does not hold once HI are introduced between the beads. While stiff dumbbells stretch slightly more due to HI, stiff chains stretch significantly less (the contrast reverses at high elasticity). Furthermore, while the distribution of extension for dumbbells continues to exhibit a power-law from its equilibrium value to its maximum value, the shape of the distribution for chains is strongly modified such that a power-law range is not clearly identifiable.
These findings suggest that simulations based on the simple dumbbell model may be missing important physics associated with HI-induced conformation-dependent drag, an issue that was raised decades ago in the context of turbulent flows by Hinch [16]. Our results therefore provide incentive for future work to incorporate chains with HI into two-way coupled EL simulations. Alternatively, one could explore using a reduced-order model in the form of a modified dumbbell that has a conformation-dependent drag [14, 16, 38]; but this raises the question of what functional dependence should be considered for the drag coefficient. One could determine a drag dependence empirically from experimental or simulation data [18, 8], or deduce it via blob theory [11]. Either way, it would be helpful to perform large-ensembles of simulations, to generate fitting data or to compare different models; the Gaussian random model will be helpful in this regard by serving as a surrogate for the turbulent flow. Consistent coarse-graining of HI-endowed chains would also require knowledge of how HI effects scale with the number of beads, in the large bead-number limit. Measuring and understanding this scaling behavior is an important task for the future.
@@
In homogeneous isotropic turbulence that is devoid of mean shear, chains with and without HI [39] do not exhibit persistent folded configurations, as demonstrated by their smooth distributions of extension. The strong influence on the behaviour of chains of the anisotropy of the turbulent flow offers an interesting counter-point to the mild influence of the non-Gaussian distribution of strain-rates (evidenced by the success of the Gaussian random flow in emulating isotropic turbulence with and without HI [39]). It appears that some statistical features of the flow—degree of isotropy—do strongly affect chain stretching, while other features—non-Gaussian extreme straining—do not. With this in mind, it would be interesting to study the dynamics of chains in an anisotropic Gaussian random flow, and to examine how the likelihood of persistent folded configurations depends on the degree of anisotropy. More generally, the possible effect of the statistical properties of the flow on the stretching of polymers should be borne in mind while developing multiscale models for turbulent polymer solutions, and models should therefore be tested in different canonical turbulent flows.
@@ Acknowledgements.
We thank Samriddhi Sankar Ray (ICTS, Bengaluru) for sharing his database of Lagrangian trajectories in homogeneous and isotropic turbulence. A.G. gratefully acknowledges the PhD Fellowship received from the IITB-Monash Research Academy (year 1) and from the Prime Ministers Research Fellowship Scheme (years 2-4). The work was supported by the Indo–French Centre for the Promotion of Advanced Scientific Research (IFCPAR/CEFIPRA, project no. 6704-1). J.R.P. acknowledges his Associateship with the International Centre for Theoretical Sciences (ICTS), Tata Institute of Fundamental Research, Bengaluru, India. D.V. acknowledges the support of Agence Nationale de la Recherche through Project No. ANR-21-CE30-0040-01. Simulations were performed on the IIT Bombay workstations Gandalf (procured through DST-SERB grant SRG/2021/001185) and Aragorn (procured through the IIT-B grant RD/0519-IRCCSH0-021).
@@ Appendix A DNS data and the random model for velocity gradients
The DNS corresponds to stationary homogeneous isotropic turbulence, with a Kolmogorov time-scale $\tau_{\eta}=3.72\times 10^{-2}$ (in simulation units). Note that $\tau_{\eta}$ is the turnover time-scale of the small dissipative eddies in the flow. Relative to this timescale, the Lyapunov exponent of the tracer trajectories (used for defining $\mathrm{Wi}$ in Sec. II.4) is found to be $\lambda=0.136/\tau_{\eta}$, in keeping with previous simulations of isotropic turbulence [63]. We have computed the Lyapunov exponent from the velocity gradient data using the continuous $QR$ method [56]. We have also calculated the Lagrangian correlation time-scales of the strain-rate and vorticity, $\tau_{S}$ and $\tau_{\varOmega}$, which will be needed to construct the random velocity gradient time series (as discussed below). Defining the rate-of-strain and rotation tensors as $\bm{S}=(\nabla\bm{u}+\nabla\bm{u}^{\top})/2$ and $\bm{\varOmega}=(\nabla\bm{u}-\nabla\bm{u}^{\top})/2$, we calculate the autocorrelation functions of $S_{11}$ and $\varOmega_{12}$ and find an approximately exponential decay in time. On integrating these functions, we obtain $\tau_{S}=2.20\,\tau_{\eta}$ and $\tau_{\varOmega}=8.89\,\tau_{\eta}$, in agreement with previous numerical simulations at comparable $R_{\lambda}$ [83]. The value of the Kolmogorov time-scale $\tau_{\eta}$, mentioned above, is determined from $S_{11}$, using isotropy, as $\tau_{\eta}=\left({15\langle S_{11}^{2}\rangle}\right)^{-1/2}$.
@@
where $\zeta_{i}(t)$ ($i=1,\dots,5$) and $\varpi_{i}(t)$ ($i=1,2,3$) are independent zero-mean unit-variance Gaussian random variables, with exponentially decaying autocorrelation functions and integral times $\tau_{S}$ and $\tau_{\varOmega}$, respectively. Consequently, $S_{ij}$ and $\varOmega_{ij}$ are Gaussian matrices that satisfy $\langle S_{ij}\rangle=\langle\varOmega_{ij}\rangle=0$,
$\langle S_{ik}(t)S_{jl}(0)\rangle$ $=3A^{2}\left(\delta_{ij}\delta_{kl}+\delta_{il}\delta_{jk}-\frac{2}{3}\delta_{ik}\delta_{jl}\right)\,e^{-t/\tau_{S}},$ (10)
@@
By setting the integral times $\tau_{S}$ and $\tau_{\varOmega}$ to match the corresponding values of the Lagrangian DNS database, we obtain a fluctuating time series $\bm{\kappa}(t)$ whose temporal correlation approximates that of the turbulent velocity gradient (the correlated random numbers $\zeta_{i}(t)$ and $\varpi_{i}(t)$ are generated using the algorithm of Fox et al. [84]). The magnitude of the random velocity gradient is set by the constant $A$, which therefore also determines the Lyapunov exponent $\lambda$. Choosing $A=2.538$ yields approximately the same value for $\lambda$ as in the DNS.

### v2

@@ V Random gradients emulate turbulence
Before concluding, we check whether the effects of HI on turbulent polymer stretching can be predicted by using random Gaussian velocity gradients, i.e., by replacing the turbulent flow with a synthetic random flow. As discussed in Sec. II.4, the time series of random gradients is constructed so as to have the same Lyapunov exponent and similar temporal correlation properties as the Lagrangian trajectories from the DNS (see Appendix A for details). If the Gaussian gradients work well, then one can study the extensional dynamics of polymer chains without having to perform a DNS. This would facilitate the development of coarse-grained models, like dumbbells with a conformation-dependent drag [18, 20, 11, 41], by allowing for the comparison of polymer models in a fluctuating flow without having to simulate the Navier-Stokes equations. Data-driven modelling approaches [80, 81, 82, 83] which require large amounts of data would particularly benefit from using the random flow model, since one can then perform very long simulations without the need for long Lagrangian trajectories from a DNS.
The PDF of extension predicted using the Gaussian gradients (GG) are compared to the results from the DNS, for twenty-bead chains without HI in Fig. 9(a). The comparison with HI is presented in Fig. 9(b) for $h^{\star}=0.2$ and in Fig. 9(c) for $h^{\star}=0.49$. The DNS and GG results agree reasonably well even in the presence of HI. Importantly, the differences between the PDFs from the DNS and the GG model are much smaller than the modifications produced by HI [compare the differences in Fig. 9(b) to those in Fig. 4(b)]. This means that we can replace the DNS by the GG random model without compromising our ability to study and understand the effects of HI on polymer stretching in fluctuating flows.
After averaging over the PDFs, the results for the mean extension $\langle R\rangle/R_{\mathrm{eq}}$ from the DNS and the GG random model match well, as shown in Fig. S4 of the SM [76]. We have also found that the effects of HI on the persistence time in stretched and coiled states (discussed in Sec. IV.3) are reproduced by the GG random model.
It is important to note that, though the principal Lyapunov exponent $\lambda$ is the same in the two flows, the distribution of $\nabla\bm{u}$ is very different. The velocity-gradient in turbulence is well-known to exhibit strong extreme-valued fluctuations [73, 74] that will be absent in the Gaussian model. Indeed, we find that the PDF of the strain-rate in the turbulent DNS has strongly non-Gaussian, flared tails and differs markedly from the corresponding PDF in the Gaussian flow (see Fig. S8 in the SM [76]). The close agreement of the extension statistics in these two flows implies, therefore, that the distribution of $\nabla\bm{u}$, particularly, the presence of extreme events, does not strongly affect the statistics of polymer stretching. As explained in Ref. [42], this lack of sensitivity to extreme straining is a consequence of the elasticity of the polymer: the polymer extension does not depend on the instantaneous strain-rate but rather effectively integrates the fluctuating strain rate over a time period that scales with the elastic relaxation time.
The relative unimportance of extreme strain-rate events has implications for how individual polymer stretching is altered by elastic feedback forces through its modification of the global flow. In this work, we have accounted for the local effect of feedback forces by including HI but neglected the global effect, which will change the distribution of the fluctuating velocity gradient experienced by the polymer. In homogeneous isotropic turbulence, elastic feedback reduces both the mean magnitude of the strain-rate and its extreme fluctuations [37, 38, 84, 50]. The results of this section suggest that the loss of extreme events will not matter and that polymer extension will be affected primarily by $\lambda$. This change can be accounted for by adjusting $\mathrm{Wi}$, and so we expect that the distribution of polymer extension in the presence of global elastic feedback will correspond closely to that for a lower $\mathrm{Wi}$ without global feedback. This has been demonstrated in two-way coupled simulations of dumbbells without HI by Ref. [50]. Our present results suggest that the same behavior will be obtained for chains with HI.
@@ VI Concluding remarks
Simulations of viscoelastic turbulence typically adopt dumbbell-based descriptions of the polymer, either indirectly through continuum constitutive equations such as Oldroyd–B or FENE-P [85], or directly through Eulerian–Lagrangian (EL) approaches in which explicit dumbbells are evolved alongside the Navier–Stokes equations [86, 87, 4, 50, 88]. A central modeling question is what is lost in such very coarse, fixed-drag descriptions when the underlying polymer is a long flexible chain for which intramolecular hydrodynamic interactions (HI) generate shielding and hence conformation-dependent hydrodynamic coupling. To gain insight into this issue—raised decades ago in the context of turbulent flows by Hinch [20]—we have examined the effect of HI on the stretching of a polymer in turbulence, considering a range of $\mathrm{Wi}$ and different levels of coarse-graining (from dumbbells to twenty bead chains).
Within the dumbbell idealization ($N_{\mathrm{b}}=2$), we find that adding explicit HI produces negligibly small changes in the stretching statistics. However, this should not be interpreted as validating fixed-drag dumbbell models as faithful surrogates for physically realistic chains with intramolecular HI. Instead, our fine-graining results show that once the polymer is allowed to form physically coiled configurations, HI induces hydrodynamic shielding and thereby slows migration between coiled and stretched states. This mechanism reorganizes the stretching dynamics in a way that cannot be reduced to a simple horizontal shift of the $\mathrm{Wi}$ axis: it steepens the coil–stretch transition and reshapes the PDF of extension at small and intermediate $R$. For a fixed nonzero value of the HI parameter $h^{\star}$, the deviation between HI-endowed and free-draining predictions is most pronounced near $\mathrm{Wi}^{\star}\simeq 1/2$, where polymers repeatedly traverse coiled and stretched configurations. At large $\mathrm{Wi}^{\star}$, where chains spend most of their time strongly stretched and inter-bead separations are large, the influence of HI on mean extension becomes comparatively weak.
These effects of HI grow systematically with $N_{\mathrm{b}}$, over the modest range of two to twenty considered in this study. We expect the influence of HI to continue to increase with further fine-graining, i.e, as $N_{\mathrm{b}}$ is increased toward the Kuhn-step resolution $N_{\textsc{k}}$, until a saturation is attained for a large number of beads. The asymptotic $N_{\mathrm{b}}$-dependence of HI-endowed chain dynamics is central to successive fine-graining schemes [52], which are well developed for equilibrium properties [7] and viscometric flows [65, 77, 52, 78, 79]. Extending such ideas to turbulence would require simulations at $N_{\mathrm{b}}\gg 20$ or the development of extrapolation strategies based on how leading-order corrections scale with $N_{\mathrm{b}}$ at fixed $N_{\textsc{k}}$. Understanding this scaling in turbulent flows is an important direction for future work.
@@
The most obvious characteristic of multibead chains, which is absent in dumbbells, is that elastic relaxation occurs on multiple time-scales. However, this feature alone does not lead to differences in the dynamics of the end-to-end extension. Indeed, the transient and stationary statistics of a Rouse chain in a homogeneous isotropic turbulent flow can be mapped to that of a dumbbell using the Jin-Collins mapping [48, 42]. Our current work has shown that this simple equivalence does not hold once HI are introduced. While stiff dumbbells stretch slightly more due to HI, stiff chains stretch less (the contrast reverses at higher elasticity). Furthermore, while the distribution of extension for dumbbells continues to exhibit a power-law, from its equilibrium value to its maximum value, the shape of the distribution for chains is strongly modified such that a power-law range is not clearly identifiable. The important implication of these results is that one cannot capture, not even qualitatively, the effects of HI on a polymer by simply adding HI to a dumbbell.
Now, the elastic dumbbell certainly is and will remain a valuable model; it enables computationally tractable simulations of viscoelastic turbulence by providing a minimal representation of the essential physics of flow-induced stretching and entropic relaxation. However, if the goal is predictive capability for engineering design—for example, to predict the extent of drag reduction achievable by a given polymer of a certain molecular weight—then our results show that the dumbbell is inadequate because it misses systematic chain-level effects of HI. This provides motivation for developing reduced-order models that remain computationally tractable yet emulate chain-level hydrodynamics, for example dumbbells endowed with an extension-dependent (or otherwise conformation-aware) drag [18, 20, 41, 11]. To calibrate such models, it would be helpful to perform large-ensembles of simulations to generate fitting data or to compare polymer models; the Gaussian random flow will be helpful in this regard by serving as a surrogate for the turbulent flow.
@@
In homogeneous isotropic turbulence that is devoid of mean shear, chains with and without HI [48, 16, 42] do not exhibit persistent folded configurations, as demonstrated by their smooth distributions of extension. The strong influence on the behaviour of chains of the anisotropy of the turbulent flow offers an interesting counter-point to the mild influence of the non-Gaussian distribution of strain-rates (evidenced by the success of the Gaussian random flow in emulating isotropic turbulence with and without HI [42]). It appears that some statistical features of the flow—degree of isotropy—do strongly affect chain stretching, while other features—non-Gaussian extreme straining—do not. With this in mind, it would be interesting to study the dynamics of chains in an anisotropic Gaussian random flow, and to examine how the likelihood of persistent folded configurations depends on the degree of anisotropy. More generally, the possible effect of the statistical properties of the flow on the stretching of polymers should be borne in mind while developing multiscale models for turbulent polymer solutions, and models should therefore be tested in different canonical turbulent flows.
@@ Acknowledgements.
We thank Samriddhi Sankar Ray (ICTS, Bengaluru) for sharing his database of Lagrangian trajectories in homogeneous and isotropic turbulence. A.G. gratefully acknowledges the PhD Fellowship received from the IITB-Monash Research Academy (year 1) and from the Prime Minister’s Research Fellowship Scheme, Ministry of Education, Govt. of India (years 2-4). The work was supported by the Indo–French Centre for the Promotion of Advanced Scientific Research (IFCPAR/CEFIPRA, project no. 6704-1). J.R.P. acknowledges his Associateship with the International Centre for Theoretical Sciences (ICTS), Tata Institute of Fundamental Research, Bengaluru, India. D.V. acknowledges the support of Agence Nationale de la Recherche through Project No. ANR-21-CE30-0040-01. Simulations were performed on the IIT Bombay workstations Gandalf (procured through DST-SERB grant SRG/2021/001185) and Aragorn (procured through the IIT-B grant RD/0519-IRCCSH0-021).
@@ Appendix A DNS data and the random model for velocity gradients
The DNS corresponds to stationary homogeneous isotropic turbulence, with a Kolmogorov time-scale $\tau_{\eta}=3.72\times 10^{-2}$ (in simulation units). Note that $\tau_{\eta}$ is the turnover time-scale of the small dissipative eddies in the flow. Relative to this timescale, the Lyapunov exponent of the tracer trajectories (used for defining $\mathrm{Wi}$ in Sec. II.3) is found to be $\lambda=(0.136\pm 0.0006)/\tau_{\eta}$. We have computed the Lyapunov exponent from the velocity gradient data using the continuous $QR$ method [67, 68, 69], of which a brief description follows. Consider a line element $\bm{\ell}(t)$ evolving along a fluid trajectory $\bm{x}(t)$. The Lyapunov exponent is defined as $\lambda=\lim_{t\to\infty}t^{-1}|\ell(t)/\ell(0)|$. In a chaotic flow, $\lambda>0$ and therefore $\ell(t)$ grows as $\ell(t)\sim\ell(0)\mathrm{e}^{\lambda t}$ as $t\to\infty$. Thus, $\lambda$ measures the asymptotic growth rate of line elements [90]. In principle, $\lambda$ could be calculated directly by solving the evolution equation for $\bm{\ell}(t)$:
$\frac{\mathrm{d}\bm{\ell}}{\mathrm{d}t}=\mathbb{A}(t)\bm{\ell}$ (9)
with $\mathbb{A}_{ij}(t)=\partial_{j}u_{i}(\bm{x}(t),t)$. However, this approach is impractical in a numerical calculation, precisely because $\ell(t)$ grows exponentially. The $QR$ method is one of the methods that have been designed to circumvent this issue. Since (9) is a system of linear ordinary differential equations, its solution has the form $\bm{\ell}(t)=\mathbb{W}(t)\bm{\ell}(0)$, where $\mathbb{W}(t)$ is the fundamental matrix of the system. Therefore, the Lyapunov exponent can also be written as $\lambda=\lim_{t\to\infty}t^{-1}|\mathbb{W}(t)\hat{\bm{\ell}}(0)|$. In the $QR$ method, $\mathbb{W}(t)$ is decomposed into an orthonormal matrix $\mathbb{Q}(t)$ and an upper triangular matrix $\mathbb{R}(t)$, with positive diagonal entries. While the entries of $\mathbb{R}(t)$ can grow in time, $\mathbb{Q}(t)$ is necessarily bounded. The continuous $QR$ algorithm solves an ordinary differential equation for $\mathbb{Q}$ along the trajectory $\bm{x}(t)$ and recovers $\lambda$ from the temporal evolution of $\mathbb{Q}(t)$ and $\mathbb{A}(t)$ alone; any manipulation of exponentially growing quantities is thus avoided (see Refs. [67, 68, 69] for further details).
We have also calculated the Lagrangian correlation time-scales of the strain-rate and vorticity, $\tau_{S}$ and $\tau_{\varOmega}$, which will be needed to construct the random velocity gradient time series (as discussed below). Defining the rate-of-strain and rotation tensors as $\bm{S}=(\nabla\bm{u}+\nabla\bm{u}^{\top})/2$ and $\bm{\varOmega}=(\nabla\bm{u}-\nabla\bm{u}^{\top})/2$, we calculate the autocorrelation functions of $S_{11}$ and $\varOmega_{12}$ and find an approximately exponential decay in time. On integrating these functions, we obtain $\tau_{S}=2.20\,\tau_{\eta}$ and $\tau_{\varOmega}=8.89\,\tau_{\eta}$, in agreement with previous numerical simulations at comparable $R_{\lambda}$ [91]. The value of the Kolmogorov time-scale $\tau_{\eta}$, mentioned above, is determined from $S_{11}$, using isotropy, as $\tau_{\eta}=\left({15\langle S_{11}^{2}\rangle}\right)^{-1/2}$.
@@
where $\zeta_{i}(t)$ ($i=1,\dots,5$) and $\varpi_{i}(t)$ ($i=1,2,3$) are independent zero-mean unit-variance Gaussian random variables, with exponentially decaying autocorrelation functions and integral times $\tau_{\mathrm{s}}$ and $\tau_{\varOmega}$, respectively. Consequently, $S_{ij}$ and $\varOmega_{ij}$ are Gaussian matrices that satisfy $\langle S_{ij}\rangle=\langle\varOmega_{ij}\rangle=0$,
$\langle S_{ik}(t)S_{jl}(0)\rangle$ $=3A^{2}\left(\delta_{ij}\delta_{kl}+\delta_{il}\delta_{jk}-\frac{2}{3}\delta_{ik}\delta_{jl}\right)\,e^{-t/\tau_{\mathrm{s}}},$ (11)
@@
By setting the integral times $\tau_{\mathrm{s}}$ and $\tau_{\varOmega}$ to match the corresponding values of the Lagrangian DNS database, we obtain a fluctuating time series $\bm{\kappa}(t)$ whose temporal correlation approximates that of the turbulent velocity gradient (the correlated random numbers $\zeta_{i}(t)$ and $\varpi_{i}(t)$ are generated using the algorithm of Fox et al. [92]). The magnitude of the random velocity gradient is set by the constant $A$, which therefore also determines the Lyapunov exponent $\lambda$. Choosing $A=2.538$ yields approximately the same value for $\lambda$ as in the DNS.